OpenStudy (anonymous):

1/4+1/6 is the lcd 24?

OpenStudy (lalaly):

12

OpenStudy (curry):

no 12 the asnwer is 5 / 12

OpenStudy (anonymous):

yeah twelve

OpenStudy (curry):

dont forget medal

OpenStudy (anonymous):

To find LCD's I like to make a chart of the factors: \[\begin{array}{|c|c|c} Denominator & 2 & 3 \\ \hline 4 & 2 & 0\\ 6 & 1 & 1 \\ \hline Max & 2 & 1\end{array}\]\[\implies LCD = 2^2 \cdot 3^1 = 4\cdot 3 = 12\]

OpenStudy (anonymous):

It's slightly overkill for this, but it helps when you have a lot of numbers for which you're trying to find the LCD.

OpenStudy (smurfy14):

12